OpenStudy (anonymous):

what is 2/y +5/2=4-2/3y i got y=12/5

OpenStudy (lgbasallote):

\[\frac 2y + \frac 52 = 4 - \frac 2{3y}\] ???

OpenStudy (anonymous):

i multiplied everything by 6y

OpenStudy (lgbasallote):

is that the question?

OpenStudy (anonymous):

just solve for y..."solve each equation" is what the book says

OpenStudy (lgbasallote):

but did i write your question right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

if you multiply everything by 6y you should get \[\implies 12 + 15y = 24y -4\] right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so iff you add 4 to both sides and subtract 15y to both sides...you should get.. \[\implies 16 = 9y\] right?

OpenStudy (anonymous):

wait i got something different

OpenStudy (anonymous):

after multiplying everything by 6y i got 12+15y=24y-4y

OpenStudy (lgbasallote):

6y/3y = 2...then 2 x 2 = 4 there should be no y

OpenStudy (lgbasallote):

you made a mistake

OpenStudy (anonymous):

okay walk me through this

OpenStudy (lgbasallote):

what dont you get?

OpenStudy (lgbasallote):

\[\huge \frac{2}{3y} \times 6y \implies \frac{2}{\cancel{3y}} \times \cancel{6y}^2 \implies 2\times 2 \implies 4\]

OpenStudy (lgbasallote):

that's the only mistake you did

OpenStudy (anonymous):